OpenStudy (anonymous):

First six shipments: 10, 16, 22, 28, 34, 40 A factory is producing a linear number of objects in each shipment. If x = the shipment number, which function can be used to determine how many objects are in that shipment?

OpenStudy (anonymous):

What is 16-10= 22-16= 28-22= 34-28= 40-34= Can you find the slope of the linear function?